a point p is at a distance of 10 from a point (2,3).find the coordinates of the point p if the y coordinate is twice the x coordinate?

Answer by stanbon(75887)

a point p is at a distance of 10 from a point (2,3).find the coordinates of the point p if the y coordinate is twice the x coordinate?
------------
Let P be (k,2k)::
Equation:
(k-2)^2 + (2k-3)^2 = 10^2
-------------------------------
K^2-4k+4 + 4k^2-12k+9 = 100
5k^2 - 16k - 87 = 0
------
k = 6.07
2k = 12.14
------------------
Cheers,
Stan H.